OpenStudy (anonymous):

The predicted equation for total sales over time is given by the equation: Total = 120*ln(t^2 + 0.63) + 50.7 where t is given in months and "Total" is given in thousands of units. Each copy of the movie sells for $14 and costs $7.7 to produce. The sunk cost of producing the film that has not been recovered from theater sales is $5100000 a.) Develop a general, linear profit equation, assuming the cost is linear and using x to represent a single unit of sales. Profit = b.) Find an equation that represents total profit over time by using a composition of functions to change from x units to time

OpenStudy (anonymous):

Profit is revenue minus cost. Let \(x\) be the quantity sold. Revenue in terms of quantity is given by: \[ r(x) = 14x \]The cost in terms of quantity sold is given by:\[ c(x) = 7.7x+5100000 \]This means profit would be\[ p(x) = r(x) - c(x) = 14x-(7.7x+5100000)=6.3x-5100000 \]

OpenStudy (anonymous):

Now, to find this in terms of time, remember that: \[ x(t) = 120\ln(t^2 + 0.63) + 50.7 \]

OpenStudy (anonymous):

Thank you Wio, however for question a, we are looking for "x" to be a single unit of sales

OpenStudy (anonymous):

Yes, which is why \(p(x)\) would be profit in terms of sales. Since \(x\) is number of sales.

OpenStudy (anonymous):

Oh, ok...

OpenStudy (anonymous):

Where as profit in terms of time is just \[\begin{split} (p\circ x)(t) &= p(120\ln(t^2 + 0.63) + 50.7)\\ &=6.3(120\ln(t^2 + 0.63) + 50.7)−5100000\\ &=756\ln(t^2+0.63)-5099680.59 \end{split}\]

OpenStudy (anonymous):

Thank you Wio. That is really helpful. So, for break even quantity would I just set 0=6.3x−5100000 and solve for x correct?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

ok great. I got 809523.809524 for break even qty. How long will it take in months to reach this quantity. How would I set up the equation?

OpenStudy (anonymous):

Still there Wio?